A thin cylindrical pipe with an inside diameter of 200 mm is 4 mm thick and fitted with end covers. It is subjected to an internal pressure of 4,8 MPa. Calculate the circumferential and longitudinal stress induced. Which axially applied compressive load would reduce the longitudinal stress to zero?\u00a0\n"}]}
1
Expert's answer
2020-12-08T02:41:35-0500

"\delta_(circum)=\frac{Pd}{2t}=\frac{4.8*200}{2*4}=120 Mpa"


"\delta_(long)=\frac{Pd}{4t}=\frac{4.8*200}{4*4}=60 Mpa"


To make longitudinal stress zero when compressive load is applied. That is to say combined stress in the longitudinal direction due to internal pressure and compressive load should be zero.


"\delta_(long)comp+\delta_(long)interpressure=0"


"60-\frac{F}{\pi*Dt}=0"


"F= 60*\pi*200*4=150.7946 kW". This is compressive load axially